OpenStudy (anonymous):

I need help please! how would I solve this for u B=d/1-u

OpenStudy (rock_mit182):

hi :) is like that ? \[B = \frac{ d }{ (1-u) }\]

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the () aren't there

OpenStudy (rock_mit182):

is better with them i guees

OpenStudy (rock_mit182):

multiply both sides of the eqution by (1-u)

OpenStudy (anonymous):

so would it be b(1-u)=d

OpenStudy (rock_mit182):

good;) now distribuite the b

OpenStudy (rock_mit182):

i mean multiply b,

OpenStudy (anonymous):

b-bu=d

OpenStudy (anonymous):

im confused

OpenStudy (rock_mit182):

now what do you think ? you are so near to solve this

OpenStudy (anonymous):

divide b

OpenStudy (rock_mit182):

YES !

OpenStudy (rock_mit182):

:)

OpenStudy (anonymous):

thanks

OpenStudy (rock_mit182):

you're welcome

OpenStudy (rock_mit182):